### v2

# Towards Optimal Aggregation of Varying Range Dependencies in Haze Removal

Abstract
Haze removal aims to restore a clear image from a hazy input. Existing methods have shown significant efficacy by capturing either short-range dependencies for local detail preservation or long-range dependencies for global context modeling. Given the complementary strengths of both approaches, a intuitive advancement is to explicitly integrate them into a unified framework. However, this potential remains underexplored in current research. In this paper, we propose DehazeMatic, which leverages the proposed Transformer-Mamba Dual Aggregation block to simultaneously and explicitly captures both short- and long-range dependencies through dual-path design for improved restoration. To ensure that dependencies at varying ranges contribute optimally to performance, we conduct extensive experiments to identify key influencing factors and determine that an effective aggregation mechanism should be guided by the joint consideration of haze density and semantic information. Building on these insights, we introduce the CLIP-enhanced Dual-path Aggregator, which utilizes the rich semantic priors encapsulated in CLIP and the estimated haze density map, derived from its powerful generalization ability, to instruct the aggregation process. Extensive experiments demonstrate that DehazeMatic outperforms sort-of-the-art methods across various benchmarks.
Figure 1: Illustration of factors influencing the relative importance between short-range and long-range dependencies. We train a Transformer-based dehazing network with a global receptive field and synthesis hazy images with different densities based on the Atmospheric Scattering Model [49]. (a) The horizontal axis represents the Euclidean distance between the current token and all other tokens, while the vertical axis represents the average attention weights between the current token and all tokens at the corresponding distance. The curve data is derived from all tokens of all images in the test set. (b) The bottom visualizes the attention map between the chosen tokens (indicated by the tails of red arrows) and all other tokens. On the left is the visualization of tokens at the same position corresponding to the same ground truth image, but with different haze densities. On the right is a visualization of tokens with different semantic information on the same homogeneous haze image. More analysis is in supplementary material.
1 Introduction
Single image dehazing aims to recover a clear image from a hazy input, serving as a crucial pre-processing step for high-level vision tasks under hazy conditions, such as object detection [39] and semantic segmentation [53].
Figure 2: Illustration of CLIP [52]’s potential to perceive haze and its density. We use the ViT-B/32 OpenCLIP [26] model, pre-trained on the LAION dataset, for our experiments. The values above the images indicate the PSNR and SSIM metrics, which reflect haze density from a quantitative perspective. The values below the images represent the similarity scores between the images and the paired prompts from CLIP. Notably, as haze density increases, the similarity score with the haze-describing prompt also increases.
Current data-driven methods can be categorized into two classes based on the receptive field size of the operators used for feature extraction, namely (i) using convolution [5, 7, 13, 37, 55, 56, 72] or window-based self-attention [35, 34, 59, 64] and (ii) using linear self-attention [51] or State Space Model (SSM) [58, 80, 73]. The former excels at capturing short-range dependencies but lacks effective global modeling capabilities [30, 62, 12]. In contrast, the latter directly models long-range dependencies while ensuring efficient computation; however, it fails to preserve local 2D inductive bias [25, 14].
Both categories of methods have achieved excellent performance, and their respective strengths can address the limitations of the other. Therefore, a natural idea is that explicitly combining short- and long-range dependencies (i.e., leveraging the strengths of both approaches) is expected to yield even more promising results. However, this integration has not been well explored. To address this gap, we will employ a dual-path design that concurrently captures fine-grained perception from nearest neighbor features and holistic contextual cues from non-local relationships. The ensuing problem is that reasonably integrating short- and long-range dependencies from these two paths is a non-trivial task, because tokens with different characteristics in the image require varying ranges of dependencies. Simple addition or concatenation often leads to suboptimal performance. More importantly, there is still no clear guidance on how to determine the relative importance between short-range and long-range dependencies for each token.
One preliminary hypothesis is that haze density serves as a determining factor. To verify this, we conduct an experiment. As shown in Figure 1(a), higher haze density increases reliance on long-range dependencies, whereas lower haze density requires more short-range dependencies. Furthermore, additional investigation reveals that tokens with varying semantic information require different ranges of dependencies, as illustrated on the right side of Figure 1(b). Based on these findings, we conclude that haze density and semantic information jointly determine the relative importance of short-range and long-range dependencies.
In this paper, building on the above insights, we propose DehazeMatic, where Mamba and Transformer are Integrated by CLIP for image dehazing. The core building block of the DehazeMatic is the proposed Transformer-Mamba Dual Aggregation (TrambaDA) block, which employs a dual-path design. In one path, we utilize window-based self-attention [45] to capture short-range dependencies. In the other path, inspired by recent advances in SSMs [18, 19, 44, 82, 42], we incorporate the Selective Scan Space State Sequential Model (S6) block [18], which maintains global information indefinitely with linear complexity. We also adopt the four-way scanning strategy to mitigate the “direction-sensitive" problem caused by the non-causal nature of visual data [44].
Next, to ensure that both short- and long-range dependencies contribute optimally to performance, a dual-path aggregator capable of adaptively learning the relative importance of each path is required. As indicated by Figure 1, the aggregator must capture both the haze density and semantic information of each token in the input image. Recent advances in CLIP [52], pre-trained on web-scale datasets, reveal that it not only encapsulates rich semantic priors (e.g., enabling strong performance in zero-shot semantic segmentation [81, 76]), but also has the potential to perceive haze and its density (as shown in Figure 2). Motivated by this, we propose the CLIP-enhanced Dual-path Aggregator (CedA). Specifically, we treat the affinity scores between the the patch-wise image embeddings from the CLIP image encoder and the text embeddings of paired prompts describing clear and hazy conditions from the CLIP text encoder as the estimated haze density map. To find the best prompts for better estimation and avoid labor-intensive prompt engineering, these paired prompts are learned. For semantic information, we reuse the the patch-wise image embeddings to provide high-level and abstract features while avoiding additional computational overhead. To supplement the low-level details, we incorporate the input features from the TrambaDA blocks and align both representations using a bidirectional cross-attention mechanism. Finally, we generate dual-path aggregation weights based on the refined semantic information and the estimated haze density map.
Figure 3: (a) The architecture of our proposed DehazeMatic. (b) The inner structure of Transformer-Mamba Dual Aggregation (TrambaDA) block. The weights generated by CLIP-enhanced Dual-path Aggregator (CedA) are used to aggregate the outputs of dual-path.
Our contributions can be summarized as follows:
  • We propose DehazeMatic, which demonstrates the superiority of explicitly and simultaneously capturing both short- and long-range dependencies while effectively aggregating dependencies across varying ranges. It achieves SOTA performance in various benchmarks.
  • For the first time, we identify the key factors required for reasonably aggregating dependencies across different ranges in image dehazing. Based on these insights, we propose the CedA module, which performs dynamic aggregation according to haze density and semantic characteristics of each token within the image.
  • The proposed CedA module further explore the potential of CLIP in image dehazing and, for the first time, demonstrate its capability in haze density estimation, all without the need for fine-tuning the encoder.
2 Related Work
2.1 Single Image Dehazing
Due to spatially variant transmission map and atmospheric light, single image dehazing is a highly ill-posed problem. Early prior-based methods [23, 16, 31, 60, 83, 6] achieved dehazing by imposing various assumptions to estimate key parameters in the Atmospheric Scattering Model (ASM) [49]. ASM provides strong theoretical support and interpretability for these methods; however, they often fail when images deviate from statistical laws.
With the rapid advancement of deep learning [33], numerous learning-based methods have been proposed, resulting in improved performance. Early methods [7, 37] employ neural networks to estimate key parameters in the ASM and subsequently restore the haze-free images. Later, ASM-independent deep networks [54, 55, 43, 40, 57, 13, 75, 50, 36, 66, 70, 59, 17, 68, 77, 8, 15, 9, 65, 69, 11] directly estimate clear images or haze residuals. DehazeFormer [59] is a representative method that achieves efficient feature extraction through window-based self-attention and several key modifications. However, its inherently limited receptive field constrains its performance potential. To achieve a global receptive field with lower computational overhead, some methods [58, 71] incorporate frequency domain features into image dehazing, offering the additional benefit of more direct learning of high-frequency information. For the same purpose, other researchers [79, 73] introduce the Mamba [18] architecture into image dehazing. However, none of these methods address how to better integrate global semantic features with local texture features.
2.2 CLIP for Low-Level Vision Tasks
Classic vision-language models, such as CLIP [52], aim to learn aligned features in the embedding space from image-text pairs using contrastive learning. Some studies have explored leveraging the rich prior knowledge encapsulated in CLIP to assist with low-level vision tasks.
In All-in-One image restoration, some researchers [47, 1, 27] use degradation embeddings from the CLIP image encoder to implicitly guide the network in making adaptive responses. In monocular depth estimation, recent studies [74, 4, 24] use CLIP to map patches of the input image to specific semantic distance tokens, which are then projected onto a quantified depth bin for estimation. In low-light enhancement, some methods [41, 48] use text-image similarity between the enhanced results and learnable prompt pairs to train the enhancement network.
Figure 4: Illustration of the CLIP-enhanced Dual-path Aggregator generating aggregation weights. To ensure computational efficiency, the TrambaDA blocks share the same aggregation weights at each stage.
Some works also introduce CLIP into image [63] and video [53] dehazing, but they all use text-image similarity between dehazed results and contrastive prompt sets as regularization to guide the restoration process. In contrast, our method further exploits the potential of CLIP by directly incorporating its latent embeddings into the main network to guide the dehazing process.
3 Method
3.1 Overall Architecture
The architecture of DehazeMatic is depicted in Figure 3. Given a hazy image, shallow features are first extracted and then passed through a symmetric encoder-decoder framework. Each stage of encoder and decoder consists of multiple Transformer-Mamba Dual Aggregation (TrambaDA) blocks, along with corresponding downsampling or upsampling operations. We also incorporate skip connections between the encoder and decoder at each level. Finally, the output of the last decoder stage is combined with the input hazy image to produce the haze-free image.
3.2 Transformer-Mamba Dual Aggregation Block
The TrambaDA block’s backbone employs a dual-path design, with one path capturing short-range dependencies and the other capturing long-range dependencies, thereby providing complementary restoration information for image tokens with distinct characteristics.
Capturing Short-range Dependencies.  We construct this path using window-based self-attention [45], which offers stronger fitting capability through dynamic weights. Specifically, given the input feature $F_{\rm{in}}\in\mathbb{R}^{H\times W\times C}$, we first divide $F_{\rm{in}}$ into non-overlapping windows with size of $M\times M$, then obtain flattened feature $F_{\mathrm{in}}^{i}\in\mathbb{R}^{M^{2}\times C}$ of each window $i$. For simplicity, we only consider the single-head attention. Performing self-attention on non-overlapping windows can be formulated as:
$\begin{split}F_{\mathrm{out\_{short}}}^{i}=&\text{Attention}(F_{\mathrm{in}}^{i}W^{Q},F_{\mathrm{in}}^{i}W^{K},F_{\mathrm{in}}^{i}W^{V}),\\ \text{Attention}&(Q,K,V)=\text{Softmax}(\frac{{QK}^{T}}{\sqrt{d_{k}}})V,\\ F_{\mathrm{out\_{short}}}=&\{F_{\mathrm{out\_{short}}}^{1},F_{\mathrm{out\_{short}}}^{2},\cdots,F_{\mathrm{out\_{short}}}^{N}\},\end{split}$ (1)
where $N=HW/M^{2}$; $W^{Q},W^{K},W^{V}\in\mathbb{R}^{C\times d_{k}}$ are projection matrices for queries, keys, and values, respectively.
Figure 5: Illustration of estimated haze density maps. The proposed method can identify the relative haze density not only in different regions of non-homogeneous haze images but also between two homogeneous haze images.
Capturing Long-range Dependencies.  We build this path using Mamba [18]’s S6 block, which preserves the global receptive field with linear complexity and incorporates hardware-aware algorithms.
Due to the “direction-sensitive” problem caused by the non-causal nature of visual data, directly applying the S6 block on the flattened image will result in suboptimal performance [44]. Following Vmamba [44], we unfold feature map along four directions, generating four independent sequences. Each feature sequence is processed through the S6 block, and all output sequences are merged into a new sequence. This process can be represented as follows:
$\begin{split}F_{\mathrm{in}}^{d}=expand&(F_{\mathrm{in}},d),d=\{1,2,3,4\},\\ \bar{F}&{}_{\mathrm{in}}^{d}=\mathrm{S6}(F_{\mathrm{in}}^{d}),\\ F_{\mathrm{out\_{long}}}=m&erge(\bar{F}_{\mathrm{in}}^{1},\bar{F}_{\mathrm{in}}^{2},\bar{F}_{\mathrm{in}}^{3},\bar{F}_{\mathrm{in}}^{4}),\end{split}$ (2)
where $expand(\cdot)$ and $merge(\cdot)$ correspond to the scan expand and scan merge operations, respectively.
3.3 CLIP-enhanced Dual-path Aggregator
The CLIP-enhanced Dual-path Aggregator (CedA) utilizes the generated pixel-level weights to adaptively aggregate dual-path outputs, ensuring that dependencies across varying ranges contribute optimally to performance:
$\begin{split}F_{out}=W_{short}F_{\mathrm{out\_{short}}}+W_{long}F_{\mathrm{out\_{long}}}.\end{split}$ (3)
$W_{short},W_{long}\in\mathbb{R}^{H\times W}$ are the aggregation weights corresponding to short- and long-range dependencies, respectively, and their generation process is shown in Figure 4.
3.3.1 Estimating Haze Density Map
Define Estimation Pipeline.  Inspired by Figure 2, we leverage CLIP to estimate the haze density map. Specifically, given a hazy image $I_{\mathrm{haze}}\in\mathbb{R}^{H\times W\times 3}$ and a paired prompts describing haze and clear conditions $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ (e.g., [’hazy image’, ’clear image’]), we encode them into latent space using a pretrained CLIP model:
$\begin{split}F_{\mathrm{image}}=&\Phi_{\mathrm{image}}(I_{\mathrm{haze}})\in\mathbb{R}^{H_{p}\times W_{p}\times d_{c}},\\ &F_{\mathrm{text}}=\Phi_{\mathrm{text}}(T)\in\mathbb{R}^{2\times d_{c}},\end{split}$ (4)
where $\Phi_{\mathrm{text}}$ is the CLIP text encoder, and $\Phi_{\mathrm{image}}$ is the CLIP image encoder without the final pooling layer, allowing the generation of a patch-wise density map. $H_{p}$ and $W_{p}$ denote the height and width of the encoded image embedding, respectively, while $d_{c}$ denotes the hidden dimension of CLIP. As shown in Figure 2, as the haze density increases, the similarity scores between $F_{\mathrm{image}}$ and the haze-describing prompt $T_{\mathrm{haze}}$ also increase. Therefore, we treat these similarity scores as the estimated density map $\mathcal{M}_{d}$:
$\begin{split}\mathcal{M}_{d}=\mathrm{Softmax}(\mathrm{sim}(F_{\mathrm{image}},F_{\mathrm{text}}))[:,:,0],\end{split}$ (5)
where $\mathrm{sim}(\cdot,\cdot)$ is the similarity calculation, and final haze density map $\mathcal{M}_{d}\in\mathbb{R}^{H_{p}\times W_{p}}$.
Table 1: Quantitative comparisons of various methods on various dehazing benchmarks. The best results and the second best results are in bold and underline, respectively.
Methods Publication RESIDE SOTS (Homogeneous Haze) Non-homogeneous Haze Overhead
Outdoor Indoor Average NH-Haze Dense-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$Param MACs
DCP [22] TPAMI-10 19.14 0.861 16.61 0.855 17.88 0.858 10.57 0.520 10.06 0.385 - -
AOD-Net [37] ICCV-17 24.14 0.920 20.51 0.816 22.33 0.868 15.40 0.569 13.14 0.414 1.76K 0.12G
GridDehazeNet [43] ICCV-19 30.86 0.982 32.16 0.984 31.51 0.983 18.33 0.667 14.96 0.533 0.96M 21.55G
FFA-Net [50] AAAI-20 33.57 0.984 36.39 0.989 34.98 0.987 19.87 0.692 16.09 0.503 4.46M 288.86G
DeHamer [20] CVPR-22 35.18 0.986 36.63 0.988 35.91 0.987 20.66 0.684 16.62 0.560 132.45M 59.25G
SRDefog [28] ACCV-22 - - - - - - 20.99 0.610 16.67 0.500 12.56M 24.18M
MAXIM-2S [61] CVPR-22 34.19 0.985 38.11 0.991 36.15 0.988 - - - - 14.10M 216.00G
SGID-PFF [5] TIP-22 30.20 0.975 38.52 0.991 34.36 0.983 - - - - 13.87M 156.67G
PMNet [70] ECCV-22 34.74 0.985 38.41 0.990 36.58 0.988 20.42 0.730 16.79 0.510 18.90M 81.13G
MB-TaylorFormer-B [51] ICCV-23 37.42 0.989 40.71 0.992 39.07 0.991 - - 16.66 0.560 2.68M 38.50G
MITNet [58] MM-23 35.18 0.988 40.23 0.992 37.71 0.990 21.26 0.712 16.97 0.606 2.73M 16.42G
DehazeFormer [59] TIP-23 34.29 0.983 38.46 0.994 36.38 0.989 20.31 0.761 16.66 0.595 4.63M 48.64G
SCANet [21] CVPR-23 - - - - - - 19.52 0.649 15.35 0.508 2.39M 258.63G
DEANet [8] TIP-24 36.03 0.989 40.20 0.993 38.12 0.991 20.84 0.801 16.73 0.602 3.65M 32.23G
UVM-Net [79] Arxiv-24 34.92 0.984 40.17 0.996 37.55 0.990 - - - - 19.25M 173.55G
OKNet [10] AAAI-24 35.45 0.992 37.59 0.994 36.52 0.993 20.29 0.800 16.85 0.620 4.42M 39.54G
DCMPNet [77] CVPR-24 36.56 0.993 42.18 0.996 39.37 0.995 - - - - 18.59M 80.42G
DehazeMatic - 38.21 0.995 41.50 0.996 39.86 0.996 21.47 0.806 17.28 0.629 4.58M 35.50G
Train Learnable Haze/Clear Prompts.  To find the best prompts for improved estimation and avoid time-consuming prompt engineering, we use learnable prompt tokens instead of manually predefined prompts to describe abstract haze and clear conditions.
The training process consists of two stages. In the first stage, we optimize the learnable paired prompts using cross-entropy loss, enabling them to preliminarily distinguish between haze and clear conditions. Given a hazy and clear image $I_{\mathrm{haze}},I_{\mathrm{clear}}\in\mathbb{R}^{H\times W\times 3}$, we define the first-stage loss $\mathcal{L}_{1}$ as:
$\begin{split}\mathcal{L}_{1}=-(y*\log(\hat{y})+(1-y)*\log(1-\hat{y})),\\ \hat{y}=\frac{e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{\mathrm{clear}}))}}{\sum_{i\in\{\mathrm{haze},\mathrm{clear}\}}e^{cos(\mathrm{Pool}(\Phi_{\mathrm{image}}(I)),\Phi_{\mathrm{text}}(T_{i}))}},\end{split}$ (6)
where $I\in\{I_{\mathrm{haze}},I_{\mathrm{clear}}\}$ and $y$ is the label of the current image, 0 is for hazy image $I_{\mathrm{haze}}$ and 1 is for clear image $I_{\mathrm{clear}}$. Since $\Phi_{\mathrm{image}}$ is the CLIP image encoder without the final pooling layer, we apply a global pooling layer $\mathrm{Pool}()$ to obtain the latent image embedding with a shape of $\mathbb{R}^{1\times d_{c}}$.
In the second stage, our goal is to achieve better prediction of haze density. The most straightforward and effective optimization approach is regression. However, there is no ground-truth density map corresponding to the hazy image in the existing dataset. To address this, we construct triplet data $\{I_{\mathrm{haze}},I_{\mathrm{clear}},I_{\mathrm{density}}\}$ based on Atmospheric Scattering Model [49]. The second-stage loss $\mathcal{L}_{2}$ is defined as:
$\mathcal{L}_{2}=\begin{cases}\alpha_{1}\mathrm{MSE}(\mathcal{M}_{d},I_{\mathrm{density}})+\alpha_{2}\mathcal{L}_{1},&y=0\\ \mathcal{L}_{1},&y=1\end{cases}$ (7)
where $\mathrm{MSE}()$ is the mean absolute error, and $\mathcal{M}_{d}$ is obtained from $I_{\mathrm{haze}}$ and the learnable paired prompts $T=[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ according to Equation 5. $\alpha_{1}$ and $\alpha_{2}$ are the weights of different loss functions. More details is in supplementary material. Training is lightweight since there is no need to fine-tune the CLIP image encoder.
Finally, we use the learned paired prompts $[T_{\mathrm{haze}},T_{\mathrm{clear}}]$ to generate the estimated patch-wise haze density map $\mathcal{M}_{d}$. Some examples are shown in Figure 5, demonstrating that our method is an effective and universal approach applicable to both homogeneous and non-homogeneous haze.
3.3.2 Obtaining Refined Semantic Map
Considering that the pretrained image encoder of CLIP is obtained through a classification pretext task, each location in the feature map before pooling captures regional semantic information [74]. Therefore, we reuse $F_{\mathrm{image}}$ from Equation 4 and regard it as semantic information. This approach not only provides high-level information but also avoids introducing additional computational overhead.
However, relying solely on $F_{\mathrm{image}}$ leads to suboptimal performance, primarily because it is an abstract representation and lacks low-level semantic details. To address this limitation, we incorporate the input features of the TrambaDA blocks, $F_{\mathrm{input}}$, to supplement $F_{\mathrm{image}}$ and introduce a bidirectional cross-attention mechanism to align their semantic information across different scales, thereby yielding a refined semantic map $\mathcal{M}_{s}$:
$\begin{split}\mathcal{M}_{s}=\mathrm{Linear}&(\text{Attention}_{\text{input}\to\text{image}}(Q_{image},K_{input},V_{input}),\\ &(\text{Attention}_{\text{image}\to\text{input}}(Q_{input},K_{image},V_{image})).\end{split}$ (8)
$Q_{i}$, $K_{i}$ and $V_{i}$ represent the query, key, and value derived from $F_{\mathrm{i}}$ after channel reduction or adaptive pooling, where $\mathrm{i}\in\{\mathrm{image},\mathrm{input}\}$. This process is shown in Figure 4.
3.3.3 Generating Dual-path Aggregation Weights
Finally, the estimated haze density map $\mathcal{M}_{d}$ and the obtained refined semantic map $\mathcal{M}_{s}$ are jointly used to generate the dual-path aggregation weights, as follows:
$\begin{split}W_{short},W_{long}&=\mathrm{Split}(\mathrm{Linear}(\mathrm{Interpolate}(\\ &\quad\quad\quad\quad\mathrm{Concat}(\mathcal{M}_{d},\mathcal{M}_{s})))).\end{split}$ (9)
Here, $\mathrm{Concat}(\cdot)$ denotes a concatenation operation, $\mathrm{Interpolate}(\cdot)$ represents an interpolation operation used to resize $\mathcal{M}_{d}$ and $\mathcal{M}_{s}$ to match the dimensions of $F_{\mathrm{out\_{short}}}$ and $F_{\mathrm{out\_{long}}}$, $\mathrm{Linear}(\cdot)$ refers to a projection layer, and $\mathrm{Split}(\cdot)$ indicates a channel-wise splitting operation. The aggregation is subsequently performed using Equation 3.
4 Experiments
Figure 6: Visual comparisons on the NH-Haze dataset [3]. The red boxes highlight the restoration effects for thick haze and texture details. Please zoom in for the best view.
Figure 7: Visual comparisons on RESIDE SOTS Indoor dataset [38]. The red boxes and green boxes show the restoration of high-frequency and low-frequency information, respectively.
4.1 Experiment Setup
Training Details. DehazeMatic is implemented with PyTorch on NVIDIA A100 GPUs. We use Adam [32] optimizer with default parameters ($\beta_{1}=0.9$, $\beta_{2}=0.99$) and a cosine annealing strategy [46] with restarts. The initial learning rate is set to $2\times 10^{-4}$, gradually decreasing to $2\times 10^{-6}$. We train the homogeneous haze dataset for 200 epochs and the non-homogeneous haze dataset for 400 epochs. The images are randomly cropped to a size of 256$\times$256 and augmented with flipping. We use L1 and perceptual loss [29] to supervise dehazing process.
Datasets. We evaluate the proposed DehazeMatic on homogeneous (i.e., RESIDE [38]) and non-homogeneous haze datasets (i.e., NH-HAZE [3] and Dense-Haze [2]). The RESIDE [38] includes two training subsets: the Indoor Training Set (ITS) with 13,990 pairs of indoor images, and the Outdoor Training Set (OTS) with 313,950 pairs of outdoor images. Models are evaluated on the corresponding subset of the Synthetic Objective Testing Set (SOTS). The NH-HAZE [3] and Dense-Haze [2] use a professional haze generator to simulate real hazy conditions, making them suitable for evaluating model performance in real-world non-homogeneous haze removal. These datasets contain 55 and 55 paired images, respectively. We use the last 5 images as the testing set and the remaining images as the training set.
4.2 Comparison with SOTA Methods
Table 1 shows the quantitative results on various benchmarks. As the parameters of the CLIP encoder are frozen, we exclude them from the overhead, which is consistent with previous work. Notably, the additional inference time introduced by CLIP is minimal, requiring only 3.8 ms per image on an NVIDIA A100 GPU.
Performance on Non-homogeneous Haze. MITNet [58] and OKNet [10] are competitive in terms of PSNR and SSIM, respectively; however, our DehazeMatic achieves the highest performance in both metrics across all datasets. Moreover, as shown in Figure 6, MITNet [58] and OKNet [10] suffer from blue artifacts and blurred details, respectively. In contrast, DehazeMatic generates haze-free images with optimal fidelity and richer structural details while maintaining the highest quantitative scores.
Figure 8: Illustration of trade-off between performance and overhead. The line chart in (a) represents the average PSNR on the SOTS-Indoor and SOTS-Outdoor datasets, while the line chart in (b) represents the average PSNR on the NH-Haze and Dense-Haze datasets. The greater the extent to which the line chart’s value exceeds that of the bar chart, the better the model’s trade-off.
Table 2: Quantitative comparisons of various methods on real-world haze dataset RTTS [38]. The best results and the second best results are in bold and underline, respectively.
No-reference Metrics FFA-Net MB-TaylorFormer-B MITNet DehazeFormer DEA-Net OKNet DCMPNet DehazaMatic
[…]
ENTROPY $\uparrow$ 7.164 6.907 7.080 7.090 7.079 7.112 7.143 7.405
Performance on Homogeneous Haze. DehazeMatic achieves the SOTA in the average performance of indoor and outdoor scenes, and especially significantly outperforms the second-best method by 0.79 dB with fewer MACs in outdoor scenes. Figure 7 provides visual comparisons on the SOTS Indoor dataset [38]. In the green and red boxed regions, DehazeMatic exhibits superior color consistency and effectively restores finer details, respectively.
Trade-off between Performance and Overhead. As shown in Figure 8, our DehazeMatic achieves the best trade-off for both types of haze, demonstrating its high efficiency.
Generalization on Real-world Haze Dataset. To evaluate generalization, we compare DehazeMatic with several competitive methods on the real-world haze dataset RTTS [38], using models trained on the RESIDE OTS dataset [38] for inference. As shown in Table 2 , our method outperforms all others across no-reference metrics, highlighting its superior generalization ability and practical applicability.
4.3 Ablation Studies
We conduct extensive ablation studies to demonstrate the effectiveness of each component. For a fair comparison, we tune the hyperparameters of variant models to ensure that their overhead matches that of DehazeMatic.
Table 3: Ablation study of the explicit coexistence of short- and long-range dependencies.
Setting SOTS-Indoor NH-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
Only Short Dependencies 38.41 0.992 20.40 0.771
Only Long Dependencies 39.78 0.994 20.56 0.779
DehazeMatic 41.50 0.996 21.47 0.806
Only Convolution 38.29 0.992 20.33 0.767
Only Linear Attention 40.06 0.994 20.65 0.780
Conv. & Linear Attention 41.41 0.996 21.36 0.803
Table 4: Ablation study of CLIP-enhanced Dual-path Aggregator.
Setting SOTS-Indoor NH-Haze
$\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$
(a) Remove CedA Addition 39.55 0.993 20.71 0.781
Concatenation 39.80 0.994 20.74 0.784
(b) Overall Design W/o Density Map 40.88 0.995 20.91 0.792
W/o Semantic Map 41.02 0.995 21.23 0.794
Transmission Map 40.97 0.995 21.27 0.797
(c) Density Map Depth Map 41.16 0.996 21.28 0.795
Predefined Prompts 40.24 0.994 20.82 0.789
(d) Sematic Map W/o High-level 41.12 0.995 21.30 0.800
W/o Low-level 41.19 0.996 21.34 0.802
DehazeMatic 41.50 0.996 21.47 0.806
Explicit Coexistence of Short- and Long-range Dependencies. We remove either the Mamba path (i.e., long-range) or the Transformer path (i.e., short-range) in the TrambaDA blocks to construct variant models, while simultaneously excluding the CedA module. The corresponding results are shown in the first three rows of Table 3.
Additionally, to verify that its superiority is not dependent on a specific feature extractor, we use convolution [78] (i.e., short-range), linear self-attention [51] (i.e., long-range) and the proposed CedA module to construct a variant model and repeat the above experiments. The corresponding quantitative results are presented in the last three rows of Table 3, demonstrating the versatility of our idea. Although linear self-attention [51] achieves better results than the adopted Mamba block, its time cost is significantly larger, so we do not adopt it. The comprehensive results in Table 3 demonstrate that the complementary short- and long-range dependencies introduced by the dual-path design, along with their reasonable aggregation, can significantly enhance image dehazing.
Figure 9: Visual comparisons of haze density maps estimated using learned prompts and manually predefined prompts.
CLIP-enhanced Dual-path Aggregator (CedA). As shown in Table 4, we conduct extensive ablation studies on the CedA to verify the effectiveness of each component.
(a) We replace the CedA with addition and concatenation operations to verify the necessity of adaptively aggregating short- and long-range dependencies in the dual paths.
(b) We then verify the importance of jointly using haze density and semantic information for aggregation guidance by removing each component individually.
(c) Next, we investigate the estimated haze density map. First, we verify the rationale for using the density map rather than other commonly used guidance signals in dehazing (i.e., the transmission map from DCP [22] and the depth map from Depth Anything [67]) to guide the aggregation. The results indicate that haze density is a more suitable influencing factor.
Furthermore, we use manually predefined prompts to estimate the haze density map to verify the effectiveness of using learnable prompts. As shown in Table 4, predefined prompts lead to a significant performance drop, even falling behind the approach that relies solely on semantic maps for aggregation guidance. In addition, as illustrated in Figure 9, predefined prompts are entirely incapable of generating valid patch-wise haze density map.
(d) Finally, we validate the necessity of each type of information in the refined semantic map by ablating either high-level semantic information (i.e., from CLIP) or low-level semantic information (i.e., input of Tramba blocks).
5 Conclusion
In this paper, we propose DehazeMatic, a framework that employs a dual-path design to explicitly and simultaneously capture complementary local fine-grained details and global contextual cues. Based on key insights from our experiments—namely that different tokens in an image require varying ranges of dependencies and that haze density and semantic information are the primary factors influencing the required range—we introduce the CLIP-enhanced Dual-path Aggregator. It leverages the rich semantic priors and powerful generalization ability encapsulated in CLIP to ensure that each dependency contributes optimally to performance while further exploring CLIP’s potential in image dehazing. Extensive experiments demonstrate that DehazeMatic outperforms SOTA methods on various benchmarks and holds promise for inspiring future research.